OpenStudy (anonymous):

Frank purchased a boat for $7,525. He made a down payment of $475. He applied for a six-year installment loan with an interest rate of 8.4% in the amount of $7,050. What is the total cost of the boat after six years?

OpenStudy (anonymous):

P(i)t=T

OpenStudy (anonymous):

P Principal i Interest t Time T Total \[7050\times.084\times6=\]

OpenStudy (anonymous):

T + P = Tc Tc Total Cost

OpenStudy (anonymous):

7050 + 3553.20 = $10,603.20 +Down Payment of 475=$11,078.20